OpenStudy (anonymous):

a circular plaza has a radius of 25 meters.if you will walk around it, how many more will it be than if you will walk across it

OpenStudy (anonymous):

walking across d1=2r=2*25....walking around d2=2*pi*25...calculate difference between d2 and d1

OpenStudy (anonymous):

do u get it

OpenStudy (anonymous):